OpenStudy (debpriya):

Find the remainder when (33333 to the power 44444 + 44444 to the power 33333) is divided by 7 . (A.1 B.6 C.2 D.3)

OpenStudy (mathmath333):

this is it right \(\large \color{black}{\begin{align} & \dfrac{33333^{44444} + 44444^{33333}}{7}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (debpriya):

Yes

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \dfrac{33333^{44444} + 44444^{33333}}{7}\hspace{.33em}\\~\\ & =\dfrac{6^{44444} + 1^{33333}}{7}\hspace{.33em}\\~\\ & =\dfrac{(-1)^{44444} + 1}{7}\hspace{.33em}\\~\\ & =\dfrac{1 + 1}{7}\hspace{.33em}\\~\\ & =\dfrac{2}{7}\hspace{.33em}\\~\\ & =2\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

btw for which exam is this question

OpenStudy (debpriya):

Can you please explain the solution ? This question is for general aptitude exams :)

OpenStudy (mathmath333):

which part u dont get

OpenStudy (debpriya):

first first part how did we get 6 +1?

OpenStudy (debpriya):

And also the -1

OpenStudy (mathmath333):

33333 divide by 7 leaves remainder 6 and 44444 divide by 7 leaves remainder 1

OpenStudy (mathmath333):

when (a-1) is divide by a then the remainder is -1 for a = natural number greater than 1

OpenStudy (debpriya):

Oh got it ! Thank you so much for your help :)